OpenStudy (anonymous):

What is the counter example for n-0=0-n

Vocaloid (vocaloid):

uh, well, n could be anything, really... let's pick n = 1 1-0 does not equal 0-1

OpenStudy (mathstudent55):

The equation you have, n - 0 = 0 - n only works for n = 0. We know that by solving it. n - 0 = 0 - n n = -n 2n = 0 n = 0 That means any value other than n = 0 is a counterexample.

Vocaloid (vocaloid):

^ right, I didn't consider n = 0